Find dy.


xy^4 + x ^ 2 y = x + 3y

1
Expert's answer
2021-11-24T17:33:12-0500

"xy^4 +x^2y = x +3y\\
\text{Differentiating implicitly, we have that}\\
y^4 + 4xy^3\frac{dy}{dx} +2xy + x^2 \frac{dy}{dx}= 1 +3\frac{dy}{dx}\\
\text{Collecting like terms, we have that}\\
\frac{dy}{dx}(4xy^3+x^2-3)= 1-2xy-y^4\\
\implies \frac{dy}{dx} = \frac{1-2xy-y^4}{4xy^3+x^2-3}"